# Counting rational maps on $\mathbb{P}^{1}$ with prescribed local conditions

### v2

@@ 4.3. Some positive densities
Thus, we find that $Z<p\leq\sqrt[2d]{\gamma_{3}}x$. The total number of choices for $Q$ is $<(2x+2)^{2d+1}<2\cdot(2x)^{2d+1}$ when $x$ is sufficiently large. We have shown that
#$\begin{split}&\#T_{>Z}(x)\\ \leq&\sum_{Z<p<\sqrt[2d]{\gamma_{3}}x}2\cdot(2x)^{2d+1}\cdot\left(\frac{4dx}{p^{2d}}+2d\right)\\ =&O\left(x^{2d+2}\sum_{Z<p}\frac{1}{p^{2d}}\right)+O\left(x^{2d+1}\sum_{Z<p<\sqrt[2d]{\gamma_{3}}x}1\right)\\ =&O\left(x^{2d+2}\left(\sum_{Z<p}\frac{1}{p^{2d}}+\frac{1}{\log x}\right)\right).\end{split}$
@@
The set $V_{\Sigma}:=\left(\bigcap_{p\in\Sigma}G_{p}\right)\cap\left(\bigcap_{p\notin\Sigma}N_{p}^{(d)}\right)$ is the affine lifting of $\mathcal{V}_{\Sigma}$. Let
@@
$T_{>Z}(\vec{r}):=V_{>Z}^{\prime}(\vec{r})\setminus S_{>Z}(\vec{r}).$
It follows from [Poo03, Lemma 5.1] that
@@
$h_{Q}(x):=\mathfrak{r}(C_{1},\dots,C_{n-1},x,C_{n+1},\dots,C_{2d+2})\in\mathbb{Z}[x],$
@@
There exists a positive constant $\Gamma$ depending only on $d$ and the $r_{i}$’s with $21\leq i\leq 2d+$ 2 and $i\neq n$ such that
@@
Thus, we find that $Z<p\leq\Gamma^{1/d}r_{n}$. The total number of choices for $Q$ is less than $2\cdot\prod_{i\neq n}(2r_{i})$ when the $r_{i}$’s are sufficiently large. We have shown that
#$\begin{split}\#B_{>Z}(\vec{r})\leq&\sum_{Z<p\leq\Gamma^{1/d}r_{n}}\left(2\cdot\prod_{i\neq n}(2r_{i})\left(\frac{2dr_{n}}{p^{d}}+d\right)\right)\\ \leq&2d\prod_{i=1}^{2d+2}(2r_{i})\sum_{Z<p}\frac{1}{p^{d}}+2d\prod_{i=1}^{2d+2}(2r_{i})\sum_{Z<p\leq\Gamma^{1/d}r_{n}}\frac{1}{2r_{n}}\\ \leq&2d\prod_{i=1}^{2d+2}(2r_{i})\sum_{Z<p}\frac{1}{p^{d}}+2d\prod_{i=1}^{2d+2}(2r_{i})\frac{\Gamma^{1/d}}{\log(\Gamma^{1/d}r_{n})},\\ \end{split}$
@@
Let $\Sigma_{1}=\{p\in\Sigma:\ p<2d\}$, $\Sigma_{2}=\Sigma\setminus\Sigma_{1}$, and $\Omega=\{p\notin\Sigma:\ p^{2}\leq\gamma_{2}\}$.
@@
5. Rational maps of degree $2$
@@
$\bar{A}_{p}$ =$=\{P\in(\mathbb{Z}/p\mathbb{Z})^{m+1}:\ R(P)\neq 0\}\ \text{and}$
$\bar{B}_{p}$ =$=\left\{P\in(\mathbb{Z}/p\mathbb{Z})^{m+1}:\ R(P)=0\ \text{and}\ \frac{\partial R}{\partial X_{i}}(P)\neq 0\ \text{for some $0\leq i\leq m$}\right\}.$
Then the set $\bar{\mathcal{S}}_{p}:=\{P\in\mathbb{P}^{m}(\mathbb{Z}/p^{2}\mathbb{Z}):\ R(P)\neq 0\bmod p^{2}\}$ has
@@
$\bar{A}_{p}^{\prime}$ =$=\{P\in(\mathbb{Z}/p\mathbb{Z})^{6}:R(P)=0\}\ \text{and}$
$\bar{B}_{p}^{\prime}$ =$=\left\{P\in(\mathbb{Z}/p\mathbb{Z})^{6}:R(P)=0\ \text{and}\ \frac{\partial R}{\partial a}(P)=\ldots=\frac{\partial R}{\partial f}(P)=0\right\}.$
@@
$2$ $24$ $18$ $\frac{11}{21}$
$3$ $432$ $192$ $\frac{10}{13}$
$5$ $12000$ $2880$ $\frac{596}{651}$
$7$ $98784$ $16128$ $\frac{782}{817}$
@@
$\bar{C}_{p}:=\{(a,b,c,d,e,f)\in(\mathbb{Z}/p\mathbb{Z})^{6}:\ ce-bf=cd-af=bd-ae=0\}\ \text{and}$
$\bar{D}_{p}:=\{(a,b,c,d,e,f)\in(\mathbb{Z}/p\mathbb{Z})^{6}:\ e^{2}-4df=2dc-be+2af=b^{2}-4ac=0\}.$
@@
By substituting $f=(be-2dc)/(2a)$ into the first equation, we get a quadratic equation in $d$. This explains why there are at most $2$ solutions $(d,f)$ of the above system. Therefore
@@
Therefore, with respect to the weak box densities, more than $32.7$% rational maps of degree $2$ with rational coefficients have squarefree, hence minimal, resultant.

### v3

@@ 4.3. Some positive densities
Thus, we find that $Z<p\leq\sqrt[2d]{\gamma_{3}}x$. The total number of choices for $Q$ is $(2x+2)^{2d+1}<2\cdot(2x)^{2d+1}$ when $x$ is sufficiently large. We have shown that
$\begin{split}&\#T_{>Z}(x)\\ \leq&\sum_{Z<p<\sqrt[2d]{\gamma_{3}}x}2\cdot(2x)^{2d+1}\cdot\left(\frac{4dx}{p^{2d}}+2d\right)\\ =&O\left(x^{2d+2}\sum_{Z<p}\frac{1}{p^{2d}}\right)+O\left(x^{2d+1}\sum_{Z<p<\sqrt[2d]{\gamma_{3}}x}1\right)\\ =&O\left(x^{2d+2}\left(\sum_{Z<p}\frac{1}{p^{2d}}+\frac{1}{\log x}\right)\right).\end{split}$
@@
The set $V_{\Sigma}\mathrel{\mathop{\ordinarycolon}}=\left(\bigcap_{p\in\Sigma}G_{p}\right)\cap\left(\bigcap_{p\notin\Sigma}N_{p}^{(d)}\right)$ is the affine lifting of $\mathcal{V}_{\Sigma}$. Let
@@
$T_{>Z}(\vec{r})\mathrel{\mathop{\ordinarycolon}}=V_{>Z}^{\prime}(\vec{r})\setminus S_{>Z}(\vec{r}).$
It follows from [Poonensquarefree, Lemma 5.1] that
@@
$h_{Q}(x)\mathrel{\mathop{\ordinarycolon}}=\mathfrak{r}(C_{1},\dots,C_{n-1},x,C_{n+1},\dots,C_{2d+2})\in\mathbb{Z}[x],$
@@
There exists a positive constant $\Gamma$ depending only on $d$ and the $r_{i}$s with $1\leq i\leq 2d+2$ and $i\neq n$ such that
@@
Thus, we find that $Z<p\leq\Gamma^{1/d}r_{n}$. The total number of choices for $Q$ is less than $2\cdot\prod_{i\neq n}(2r_{i})$ when the $r_{i}$s are sufficiently large. We have shown that
[…]
@@
Let $\Sigma_{1}=\{p\in\Sigma\mathrel{\mathop{\ordinarycolon}}\ p<2d\}$, $\Sigma_{2}=\Sigma\setminus\Sigma_{1}$, and $\Omega=\{p\notin\Sigma\mathrel{\mathop{\ordinarycolon}}\ p^{2}\leq\gamma_{2}\}$.
@@
5. Rational maps of degree 22
@@
$\bar{A}_{p}$ $=\{P\in(\mathbb{Z}/p\mathbb{Z})^{m+1}\mathrel{\mathop{\ordinarycolon}}\ R(P)\neq 0\}\ \text{and}$
$\bar{B}_{p}$ $=\left\{P\in(\mathbb{Z}/p\mathbb{Z})^{m+1}\mathrel{\mathop{\ordinarycolon}}\ R(P)=0\ \text{and}\ \frac{\partial R}{\partial X_{i}}(P)\neq 0\ \text{for some $0\leq i\leq m$}\right\}.$
Then the set $\bar{\mathcal{S}}_{p}\mathrel{\mathop{\ordinarycolon}}=\{P\in\mathbb{P}^{m}(\mathbb{Z}/p^{2}\mathbb{Z})\mathrel{\mathop{\ordinarycolon}}\ R(P)\neq 0\bmod p^{2}\}$ has
@@
$\bar{A}_{p}^{\prime}$ $=\{P\in(\mathbb{Z}/p\mathbb{Z})^{6}\mathrel{\mathop{\ordinarycolon}}R(P)=0\}\ \text{and}$
$\bar{B}_{p}^{\prime}$ $=\left\{P\in(\mathbb{Z}/p\mathbb{Z})^{6}\mathrel{\mathop{\ordinarycolon}}R(P)=0\ \text{and}\ \frac{\partial R}{\partial a}(P)=\ldots=\frac{\partial R}{\partial f}(P)=0\right\}.$
@@
22 $24$ $18$ $\frac{11}{21}$
33 $432$ $192$ $\frac{10}{13}$
55 $12000$ $2880$ $\frac{596}{651}$
77 $98784$ $16128$ $\frac{782}{817}$
@@
$\bar{C}_{p}\mathrel{\mathop{\ordinarycolon}}=\{(a,b,c,d,e,f)\in(\mathbb{Z}/p\mathbb{Z})^{6}\mathrel{\mathop{\ordinarycolon}}\ ce-bf=cd-af=bd-ae=0\}\ \text{and}$
$\bar{D}_{p}\mathrel{\mathop{\ordinarycolon}}=\{(a,b,c,d,e,f)\in(\mathbb{Z}/p\mathbb{Z})^{6}\mathrel{\mathop{\ordinarycolon}}\ e^{2}-4df=2dc-be+2af=b^{2}-4ac=0\}.$
@@
By substituting f=(be2dc)/(2a)f=(be-2dc)/(2a) into the first equation, we get a quadratic equation in $d$. This explains why there are at most $2$ solutions $(d,f)$ of the above system. Therefore
@@
Therefore, with respect to the weak box densities, more than 32.732.7% rational maps of degree $2$ with rational coefficients have squarefree, hence minimal, resultant.